40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 6 21:11:45 2012, yuta, Update, Green Locking, improved Y arm green alignment beatygreenpower20120606.png
    Reply  Thu Jun 7 02:25:27 2012, yuta, Update, Green Locking, improved Y arm green alignment - even more beatygreenpower20120607.png
       Reply  Thu Jun 7 07:27:27 2012, Koji, Update, Green Locking, improved Y arm green alignment - even more freq_noise_AUX.pdf
Message ID: 6776     Entry time: Thu Jun 7 02:25:27 2012     In reply to: 6771     Reply to this: 6780
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: improved Y arm green alignment - even more 

[Koji, Yuta]

Summary:
  We improved the Y arm green transmission to the PSL table. It is now 197 uW.
  The improvement was done mainly by adjusting the Y arm green servo gain.

What we did:
  1. Fine-adjusted steering mirrors after the faraday on Y end table by monitoring Y arm green transmission (used Thorlabs PDA36A as a PD, C1:GCV-GREEN_TRY as a channel). We decided which way to adjust the mirrors by just pushing/pulling its mount.

  2. The output of the reflection PD on the oscilloscope seemed like the Y end frequency servo was oscillating. So, we reduced the amplitude of the frequency modulation from 2.83 V to 0.13 V.

  3. We noticed there were two TEM00, one is brighter and the other is dim. We thought this came from a mode-hopping or something. So, we changed the Y end laser temperature from 34.68 deg C to 34.13 deg C (measured). This reduced dim TEM00 and the main one got brighter. C1:GCY-SLOW_SERVO2_OFFSET was changed from 29425 to 29845.

  4. Fine-adjusted the position of the mode-matching lens by reduing LG modes.

Current green power:
  Current measured green power values are as follows.

beatygreenpower20120607.png

  Calculated value for the Y arm green transmission is ~ 600 uW, but we think we are almost at the maximum we can get. So, we have about 70% loss from the Y end table to the PSL table. There may be large loss in windows. The beam shape of the transmitted beam seems OK, but there may be some clipping.

To do:
  - Fine tune the Y end frequency servo loop. Reducing the amplitude of the frequency modulation for reducing the gain is not a very good idea.

ELOG V3.1.3-